# Evaluating Language Models as Synthetic Data Generators

### v1

@@ 1 Introduction
Post-training language models (LMs) on synthetic data is a promising approach for improving their ability to solve a wide range of tasks (Wang et al., 2023; Honovich et al., 2023; Taori et al., 2023; Liu et al., 2024b). While acquiring data through manual annotation continues to play an important role, synthetic data generation offers a scalable complement to human labeling (Viswanathan et al., 2023; Kim et al., 2023b). Hence, numerous works have proposed novel methods to effectively generate high-quality synthetic data (Xu et al., 2024a; Gunasekar et al., 2023; Yue et al., 2023, 2024).
@@
Lastly, we conduct analysis experiments for effective data generation. For instance, we find that the output format of synthetic data significantly impacts performance: data generated using JSON format shows 4.45% lower performance on average across six settings compared to free-form generation. Additionally, in budget-constrained setting, generating more data with a weaker model can outperform generating less data with stronger model. We find that generating 50K instances with GPT-4o-mini, while 3.4 times cheaper, achieves better performance than generating 10K instances with GPT-4o in two out of three settings.
@@ 2 Preliminaries: Measuring Data Generation Capabilities of LMs
Data Generation Methods.
  • Instance Generation: Given a small seed dataset ${D}_{seed}=\{(I_{i},R_{i})\mid i=1,\dots,m\}$, a few instances are randomly sampled from ${D}_{seed}$ and used as in-context demonstrations, resulting in the generation of new instances (Honovich et al., 2023; Wang et al., 2023). This process is performed iteratively until $D_{G}=\{(I_{i},R_{i})\mid i=1,\dots,n\}$ is constructed where $m<<n$. Note that the generated instances could also optionally be used as demonstrations as well.
@@
When training the student model ($S_{\text{\O}}$), we employ supervised fine-tuning (SFT), computing the loss only on response tokens. We directly use the generated data $D_{G}$ without filtering and do not consider other post-training methods, as our goal is to evaluate the raw data generation capabilities of an LM ($G$) in the most straightforward setting, not to maximize ${S}_{D_{G}}$’s benchmark performance. The hyper-parameters for training are detailed in Appendix D.
@@ 3 Experimental Setting of AgoraBench
Then, the student model is trained using data from a single domain to isolate the effect of generated data quality, as cross-domain training could introduce confounding factors through positive or negative transfer (e.g., training on code data improve math (Dong et al., 2023; Zhang et al., 2024)).
@@
  • For quality enhancement, we test scenarios where complete instances of instructions and responses already exist, but their quality needs improvement before being used for post-training - either because the instructions are too simple or the responses are not sufficiently detailed. We sample 10K instances from WebInstruct (Q-A pairs from the web requiring refinement; see Yue et al. (2024)) for instruction following and math domains. Note that WebInstruct does not contain domain labels, hence we prompt GPT-4o-mini-2024-07-18 to prepare a separate $D_{seed}$ (further details are in Appendix B). For the code domain, we use CoNaLa, which contains simple instructions paired with 1-3 line code snippets from StackOverflow (Yin et al., 2018).
@@ 4 Experimental Results of AgoraBench
Table 3: Comparison of API costs, problem-solving ability, and data generation ability: Our findings reveal that neither the strength nor the cost of an LM guarantees its effectiveness as a data generator. Note that the Llama models are instruction-tuned versions, the specific results of the LMs on each benchmark (averaged as ‘Problem Solving average’) is in Appendix C, and the AgoraBench results are averaged from Table 2.
@@ 5 What makes an effective data generator?
In the previous section, we observed an unexpected finding: LMs with weaker problem-solving ability sometimes outperform stronger LMs when generating the same amount of synthetic data under identical conditions. For a better understanding of this phenomenon, we first examine whether there exists a strong correlation between problem-solving ability and data generation ability (Section 5.1). Then we investigate whether we can predict the degree of improvement in the performance of the student model by analyzing the data generated by each LM (Section 5.2).
@@ 5.2 Can we predict the student model’s improvement by looking into the data?
  • Response Quality : We measure the quality of $R_{i}$ given $I_{i}$. We use two methods:
    • LLM-as-a-Judge Score: We prompt an LM to return a discrete score between 1 and 5 that represents the quality of $R_{i}$. We employ two LM judges: (1) Prometheus-2-8x7B (Kim et al., 2024), an open-source LM specialized on assessing LM output and (2) GPT-4o, a proprietary LM widely used as a judge. We use different score rubrics for each domain, listed in Appendix G.
    • Reward Model Score: We use a reward model to predict a scalar value score that represents the quality of $R_{i}$. We use Skywork-Reward-Llama-3.1-8B (Liu et al., 2024a), one of the top-performing reward models on Reward Bench (Lambert et al., 2024).
  • Instruction Complexity (LLM-as-a-Judge Score): We measure the difficulty of $I_{i}$ by prompting an LM to return a discrete score between 1 and 5 that represents the complexity of $I_{i}$. Similarly to evaluating response quality, we use Prometheus-2-8x7B and GPT-4o as a judge. Note that the score rubric differs compared to that for evaluating response quality and we use a different score rubrics for each domain, listed in Appendix G.
@@
Due to page limits, the full results of the intrinsic evaluation are further detailed in Appendix E.
@@
Table 4: Mean Contributions of Intrinsic Metrics to Principal Components: Each loading strength represents the average magnitude of a feature’s loadings across all principal components and the contribution are normalized values to represent the relative percentage of each feature’s loading strength in the overall component structure. ‘I.D.’ refers to metrics measuring instruction difficulty and ‘R.Q.’ refers to metrics measuring response quality. All intrinsic evaluation metrics show substantial contributions (0.189-0.256) to the principal components.
@@
Lastly, we predict data generation capabilities by performing linear regression on the top-5 principal components, weighting each component by its corresponding regression coefficient, as shown in Figure 6. Compared to using problem-solving scores alone (Figure 4), this approach yields a statistically significant relationship ($p<0.001$) with improved explanatory power ($R^{2}=0.325$). However, the moderate $R^{2}$ value suggests that additional intrinsic measurements beyond our current set might be needed to better predict data generation capabilities. We leave further exploration of this issue to future works.
6 Further Analysis Experiments
In this section, we further examine, two critical questions regarding data generation: (1) Should we prioritize quantity using cheaper LMs, or quality using more expensive ones? (Section 6.1) And (2) What is the impact of meta-prompt design, particularly when comparing structured JSON format generation against traditional free-form approaches? (Section 6.2).
6.1 Quantity or quality?
@@
6.2 Effect of Meta-prompts
@@
7 Conclusion
In this paper, we introduce AgoraBench, a benchmark that systematically evaluates LMs’ data generation capabilities through standardized settings and metrics. Our analysis reveals that models show distinct strengths across different generation methods and domains, highlighting the importance of careful data generator selection. While an LM’s data generation ability cannot be predicted solely by its cost or problem-solving abilities, we identify interpretable low-dimension principal components from intrinsic evaluation measurements that explain up to 93.4% of the variance and serve as better predictors.
Looking ahead, we envision AgoraBench enabling two key advances in the field. First, since our findings suggest that problem-solving ability is not the primary determinant of data generation quality, researchers can use our benchmark to identify the core capabilities that make an effective data generator and potentially develop specialized LMs specialized in data generation. Second, AgoraBench can serve as a practical evaluation framework for practitioners to assess and improve their data generation pipelines - they can use their custom data generation methods, seed datasets, or meta-prompts and compare against our baseline settings. Furthermore, they can leverage our systematic evaluation methodology to optimize their generation parameters before deploying large-scale data creation. Through these complementary research and applied directions, AgoraBench aims to accelerate both our theoretical understanding of language models as data generators and their practical deployment in real-world applications.
Appendix A Related Work
Conventionally, training LMs on human-crafted data was considered the de facto standard for improving an LM’s performance on downstream tasks (Mishra et al., 2022; Wei et al., 2021; Wang et al., 2022; Longpre et al., 2023). Yet, based on the in-context learning abilities of LMs (Brown et al., 2020), a series of works demonstrated that LMs could generate novel instances that could be used as post-training data (Wang et al., 2023; Honovich et al., 2022; Kim et al., 2023a).
Appendix B Response Generation Seed Dataset Construction
@@
Appendix C Problem Solving Abilities of LMs evaluated as Data Generators
The evaluation results of GPT-4o-2024-08-06, GPT-4o-mini-2024-07-18, Claude-3.5-Sonnet-2024-06-20, Llama-3.1-405B-Instruct, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct are listed in Table 7. We use the settings listed in Appendix D.
@@ Appendix C Problem Solving Abilities of LMs evaluated as Data Generators
Appendix D Details for Training and Evaluating Student Models
@@
Appendix E Intrinsic Evaluation of AgoraBench
@@
Appendix F AgoraBench Meta-prompts
@@
Appendix G Prompt for Intrinsic Evaluation

### v2

@@ 1 Introduction
Post-training language models (LMs) on synthetic data is a promising approach for improving their ability to solve a wide range of tasks (Wang et al., 2023; Honovich et al., 2023a; Taori et al., 2023; Liu et al., 2024b). While acquiring data through manual annotation continues to play an important role, synthetic data generation offers a scalable complement to human labeling (Viswanathan et al., 2023; Kim et al., 2023b). Hence, numerous works have proposed novel methods to effectively generate high-quality synthetic data (Xu et al., 2024a; Gunasekar et al., 2023; Yue et al., 2023, 2024).
@@ 2 Preliminaries: Measuring Data Generation Capabilities of LMs
Data Generation Methods.
  • Instance Generation: Given a small seed dataset ${D}_{seed}=\{(I_{i},R_{i})\mid i=1,\dots,m\}$, a few instances are randomly sampled from ${D}_{seed}$ and used as in-context demonstrations, resulting in the generation of new instances (Honovich et al., 2023a; Wang et al., 2023). This process is performed iteratively until $D_{G}=\{(I_{i},R_{i})\mid i=1,\dots,n\}$ is constructed where $m<<n$. Note that the generated instances could also optionally be used as demonstrations as well.
@@
PGR vs Intrinsic Metrics.
Note that prior works have mostly employed intrinsic metrics to measure data quality, such as response quality, instruction difficulty, and instance diversity. However, these metrics do not directly measure their impact on the ultimate goal for data generation: improving the student model. By introducing PGR (an extrinsic metric), we provide a way to evaluate this improvement directly across data generators. Without this direct measurement, it becomes challenging to assess how much of a student model’s improvement stems from the data generator itself versus other contributing factors. Hence, this makes PGR an important complement to intrinsic metrics rather than a replacement for them. We further compare the relationship of how we could use intrinsic metrics to estimate the PGR values in Section 5.2.
@@
When training the student model ($S_{\text{\O }}$), we employ supervised fine-tuning (SFT), computing the loss only on response tokens. We directly use the generated data $D_{G}$ without filtering and do not consider other post-training methods, as our goal is to evaluate the raw data generation capabilities of an LM ($G$) in the most straightforward setting, not to maximize ${S}_{D_{G}}$’s benchmark performance. The hyper-parameters for training are detailed in Appendix E.
@@ 3 Experimental Setting of AgoraBench
Then, the student model is trained using data from a single domain to isolate the effect of generated data quality, as cross-domain training could introduce confounding factors through positive or negative transfer (e.g., training on code data improve math (Dong et al., 2024; Zhang et al., 2025)).
@@
  • For quality enhancement, we test scenarios where complete instances of instructions and responses already exist, but their quality needs improvement before being used for post-training - either because the instructions are too simple or the responses are not sufficiently detailed. We sample 10K instances from WebInstruct (Q-A pairs from the web requiring refinement; see Yue et al. (2024)) for instruction following and math domains. Note that WebInstruct does not contain domain labels, hence we prompt GPT-4o-mini-2024-07-18 to prepare a separate $D_{seed}$ (further details are in Appendix C). For the code domain, we use CoNaLa, which contains simple instructions paired with 1-3 line code snippets from StackOverflow (Yin et al., 2018).
@@ 4 Experimental Results of AgoraBench
Table 3: Comparison of API costs, problem-solving ability, and data generation ability: Our findings reveal that neither the strength nor the cost of an LM guarantees its effectiveness as a data generator. Note that the Llama models are instruction-tuned versions, the specific results of the LMs on each benchmark (averaged as ‘Problem Solving average’) is in Appendix D, and the AgoraBench results are averaged from Table 2.
@@ 5 What makes an effective data generator?
In the previous section, we observed an unexpected finding: LMs with weaker problem-solving ability sometimes outperform stronger LMs when generating the same amount of synthetic data under identical conditions. For a better understanding, we examine whether there exists a strong correlation between problem-solving ability and data generation ability (Section 5.1). Then we investigate whether we can predict the degree of improvement in the performance of the student model by analyzing the data generated by each LM (Section 5.2).
@@ 5.2 Can we predict the student model’s improvement by looking into the data?
  • Response Quality : We measure the quality of $R_{i}$ given $I_{i}$. First, we use LLM-as-a-Judge where we prompt an LM to return a discrete score between 1 and 5 that represents the quality of $R_{i}$. We employ two LM judges: (1) Prometheus-2-8x7B (Kim et al., 2024), an open-source LM specialized on assessing LM output and (2) GPT-4o, a proprietary LM widely used as a judge. We use different score rubrics for each domain, listed in Appendix H. Next, we use Reward Models that predicts a scalar value score that represents the quality of $R_{i}$. We use Skywork-Reward-Llama-3.1-8B (Liu et al., 2024a), one of the top performing reward models on Reward Bench (Lambert et al., 2024).
  • Instruction Complexity (LLM-as-a-Judge Score): We measure the difficulty of $I_{i}$ by prompting an LM to return a discrete score between 1 and 5 that represents the complexity of $I_{i}$. Similarly to evaluating response quality, we use Prometheus-2-8x7B and GPT-4o as a judge. The score rubric differs compared to that for evaluating response quality and we use a different score rubrics for each domain, listed in Appendix H.
@@
Due to page limits, the full results of the intrinsic evaluation are further detailed in Appendix F.
@@
Table 4: Mean Contributions of Intrinsic Metrics to Principal Components: Each loading strength represents the average magnitude of a feature’s loadings across all principal components and the contribution are normalized values to represent the relative percentage of each feature’s loading strength in the overall component structure. ‘I.D.’ refers to instruction difficulty metrics and ‘R.Q.’ refers to response quality metrics.
@@
Lastly, we predict data generation capabilities by performing linear regression on the top-5 principal components, weighting each component by its corresponding regression coefficient, as shown in Figure 6. Compared to using problem-solving scores alone (Figure 4), this approach yields a statistically significant relationship ($p<0.001$) with improved explanatory power ($R^{2}=0.325$). However, the moderate $R^{2}$ value suggests that additional intrinsic metrics beyond our current set might be needed to better predict data generation capabilities.
6 Conclusion
In this paper, we introduce AgoraBench, a benchmark that systematically evaluates LMs’ data generation capabilities through standardized settings and metrics. Looking ahead, we envision AgoraBench enabling two key advances in the field. First, since our findings suggest that problem-solving ability is not the primary determinant of data generation quality, researchers can use our benchmark to identify the core capabilities that make an effective data generator and potentially develop specialized LMs specialized in data generation. Second, AgoraBench can serve as a practical evaluation framework for practitioners to assess and improve their data generation pipelines - they can use their custom data generation methods, seed datasets, or meta-prompts and compare against our baseline settings. Through these complementary research and applied directions, AgoraBench aims to accelerate both our theoretical understanding of language models as data generators and their practical deployment in real-world applications.
Limitations & Potential Risks
Due to compute constraints and expensive API costs, the experimental setting of AgoraBench does not cover the scenarios where (1) different base models beyond Llama-3.1-8B are used and (2) whether our findings will hold when generating more than 10K instances.
Regarding the first limitation, while Llama-3.1-8B is widely adopted as a base model in the community, making our findings particularly relevant for researchers working with similar architectures, investigating the generalizability of our results across different base models remains an important direction for future research. Of particular interest is the variation in response perplexity—one of our intrinsic evaluation metric—across different model architectures, as this metric’s behavior is inherently dependent on the base model.
For the second limitation, we conducted preliminary scaling experiments with more cost-effective language models (GPT-4o-mini, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct) generating up to 50K instances, as detailed in Appendix B. Future work could explore whether the relative effectiveness rankings among different data generators remain consistent at larger scales, and investigate potential scaling laws in relation to our proposed Performance Gap Recovered metric. Such analysis could provide valuable insights into the optimal volume of synthetic data required for specific applications and model architectures.
Synthetic data holds significant potential in aligning LMs towards human preferences and controlling their behavior. While our study primarily focuses on improvements in mathematical reasoning, coding capabilities, and instruction following abilities, the broader implications for model alignment and control warrant careful consideration. Future work could explore these aspects more explicitly, particularly investigating how synthetic data generation techniques might influence model behavior beyond task performance, including potential biases, safety considerations, and the robustness of aligned behaviors across different contexts and applications. Lastly, using a proprietary model as a data generator raises concerns about potential biases, lack of transparency in the generation process, and the need to ensure proper permissions for its use. These factors go beyond performance and could be critical for practitioners to consider when choosing a data generation approach. Addressing such issues is important to ensure ethical and responsible usage.
Acknowledgements
This research was supported in part by a gift from NEC Laboratories Europe. We thank the members of Neulab and L3Lab at CMU for helpful discussions. SW thanks Convergent Research.
Appendix A Related Work
Conventionally, training LMs on human-crafted data was considered the de facto standard for improving an LM’s performance on downstream tasks (Mishra et al., 2022; Wei et al., 2021; Wang et al., 2022; Longpre et al., 2023). Yet, based on the in-context learning abilities of LMs (Brown et al., 2020), a series of works demonstrated that LMs could generate novel instances that could be used as post-training data (Wang et al., 2023; Honovich et al., 2023b; Kim et al., 2023a).
@@ Appendix A Related Work
Appendix B Further Analysis Experiments
In this section, we further examine, two critical questions regarding data generation: (1) Should we prioritize quantity using cheaper LMs, or quality using more expensive ones? (Section B.1) And (2) What is the impact of meta-prompt design, particularly when comparing structured JSON format generation against traditional free-form approaches? (Section B.2).
B.1 Quantity or quality?
@@
B.2 Effect of Meta-prompts
@@
Appendix C Response Generation Seed Dataset Construction
@@
Appendix D Problem Solving Abilities of LMs evaluated as Data Generators
The evaluation results of GPT-4o-2024-08-06, GPT-4o-mini-2024-07-18, Claude-3.5-Sonnet-2024-06-20, Llama-3.1-405B-Instruct, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct are listed in Table 7. We use the settings listed in Appendix E.
@@ Appendix D Problem Solving Abilities of LMs evaluated as Data Generators
Appendix E Details for Training and Evaluating Student Models
@@
Appendix F Intrinsic Evaluation of AgoraBench
@@
Appendix G AgoraBench Meta-prompts
@@
Appendix H Prompt for Intrinsic Evaluation